4 times the square root of 18 times 2 times the square root of 48
Answer by richwmiller(17219)

4sqrt(18)*sqrt(48)
sqrt(48)=sqrt(4*4*3)=4sqrt(3)
sqrt(18)=sqrt(9*2)=3sqrt2
so we have 4*12sqrt(6)=48sqrt(6)